Find an equation in slope-intercept form(where possible)
Through (1/2, 5/3) and (3, 1/6)

Answer by AnlytcPhil(1116)

Find an equation in slope-intercept form(where possible) 
Through (1/2, 5/3) and (3, 1/6)

You need two formulas and knowledge of how to solve for y.

The first formula you need is the slope formula:

    y2 - y1
m = ———————
    x2 - x1

(x1, y1) = (1/2, 5/3) and (x2, y2) = (3, 1/6), so

    (1/6) - (5/3)
m = —————————————
     (3) - (1/2)
 
    (1/6) - (5/3)   (1/6) - (10/6)   -9/6    -3/2      3
m = ————————————— = —————————————— = ————— = ———— = - ———
     (3) - (1/2)     (6/2) - (1/2)    5/2     5/2      5

The other formula you need is the point-slope formula:

   y - y1 = m(x - x1)

y - (5/3) = (-3/5)[x - (1/2)]

  y - 5/3 = (-3/5)x + (3/5)·(1/2)

  y - 5/3 = (-3/5)x + 3/10

Now solve for y:

        y = (-3/5)x + 3/10 + 5/3

        y = (-3/5)x + 9/30 + 50/30

        y = (-3/5)x + 59/30           

Edwin
AnlytcPhil@aol.com